OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

4x-4=2x+4+6 4x-2x-4-4-6=0 2x-14=0 2x=14 x=7

mathslover (mathslover):

@Pinkiepie try to use this property : \(a(b-c) =ab- ac\) Can you expand 4(x-1) now?

OpenStudy (anonymous):

Thank you

mathslover (mathslover):

Hello rabab, welcome to OpenStudy . Please don't post direct solutions/answer here. We are here to help the users not to misguide them. So please, guide them step by step rather than posting the solution in a single post. You can also view , OpenStudy code-of-conduct : http://openstudy.com/code-of-conduct.

mathslover (mathslover):

^ http://openstudy.com/code-of-conduct